### v1

# Every nonsymmetric $4$-class association scheme can be generated by a digraph

Abstract.
A (di)graph $\Gamma$ generates a commutative association scheme $\mathfrak{X}$ if and only if the adjacency matrix of $\Gamma$ generates the Bose-Mesner algebra of $\mathfrak{X}$. In [17, Theorem 1.1], Monzillo and Penjić proved that, except for amorphic symmetric association schemes, every $3$-class association scheme can be generated by an adjacency matrix of a (di)graph. In this paper, we characterize when a commutative association scheme with exactly one pair of nonsymmetric relations can be generated by a digraph under the certain assumptions. As an application, we show that each nonsymmetric $4$-class association scheme can be generated by a digraph.
Key words and phrases:
Bose-Mesner algebra; association scheme generated by a digraph; nonsymmetric $4$-class association scheme.
2010 Mathematics Subject Classification:
05E30, 05C75, 05C50
1. Introduction
A $d$-class association scheme $\mathfrak{X}$ is a pair $(X,\{R_{i}\}_{i=0}^{d})$, where $X$ is a finite set, and each $R_{i}$ is a nonempty subset of $X\times X$ satisfying the following axioms (see [1, 2, 31, 32] for a background of the theory of association schemes):
  1. (i)
    $R_{0}=\{(x,x)\mid x\in X\}$ is the diagonal relation;
  2. (ii)
    $X\times X=R_{0}\cup R_{1}\cup\cdots\cup R_{d}$, $R_{i}\cap R_{j}=\emptyset~{}(i\neq j)$;
  3. (iii)
    for each $i$, $R_{i}^{\top}=R_{i^{\prime}}$ for some $0\leq i^{\prime}\leq d$, where $R_{i}^{\top}=\{(y,x)\mid(x,y)\in R_{i}\}$;
  4. (iv)
    for all $i,j,l$, the number of $z\in X$ such that $(x,z)\in R_{i}$ and $(z,y)\in R_{j}$ is constant whenever $(x,y)\in R_{l}$. This constant is denoted by $p_{i,j}^{l}$.
A $d$-class association scheme is also called an association scheme with $d$ classes (or even simply a scheme). The integers $p_{i,j}^{l}$ are called the intersection numbers of $\mathfrak{X}$. We say that $\mathfrak{X}$ is commutative if $p_{i,j}^{l}=p_{j,i}^{l}$ for all $i,j,l$. The subsets $R_{i}$ are called the relations of $\mathfrak{X}$. For each $i$, the integer $k_{i}:=p_{i,i^{\prime}}^{0}$ is called the valency of $R_{i}$. A relation $R_{i}$ is called symmetric if $i=i^{\prime}$, and nonsymmetric otherwise. An association scheme is called symmetric if all relations are symmetric, and nonsymmetric otherwise. An association scheme is called skew-symmetric if the diagonal relation is the only symmetric relation.
Let $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d})$ be a commutative association scheme. The adjacency matrix $A_{i}$ of $R_{i}$ is the $|X|\times|X|$ matrix whose $(x,y)$-entry is $1$ if $(x,y)\in R_{i}$, and 00 otherwise. By the adjacency or Bose-Mesner algebra $\mathcal{M}$ of $\mathfrak{X}$ we mean the algebra generated by $A_{0},A_{1},\ldots,A_{d}$ over the complex field. Axioms (i)(iv) are equivalent to the following:
$A_{0}=I,\quad\sum_{i=0}^{d}A_{i}=J,\quad A_{i}^{\top}=A_{i^{\prime}},\quad A_{i}A_{j}=\sum_{l=0}^{d}p_{i,j}^{l}A_{l},$
where $I$ and $J$ are the identity and all-one matrices of order $|X|$, respectively. By [17], $\mathcal{M}$ is a monogenic algebra, that is, there always exists a matrix $A\in{\rm Mat}_{X}(\mathbb{C})$ which generates $\mathcal{M}$, i.e., $\mathcal{M}=(\langle A\rangle,+,\cdot)$. We say that a matrix $A$ generates $\mathcal{M}$ if every element in $\mathcal{M}$ can be written as a polynomial in $A$.
One way to construct new association schemes is by merging or splitting relations in an existing scheme. More precisely, a partition $\Lambda_{0},\Lambda_{1},\ldots,\Lambda_{e}$ of the index set $\{0,1,\ldots,d\}$ is admissible [14] if $\Lambda_{0}=\{0\},\Lambda_{i}\neq\emptyset$ and $\Lambda_{i}^{\prime}=\Lambda_{j}$ for some $j\ (1\leq i,j\leq e)$, where $\Lambda_{i}^{\prime}=\{\alpha^{\prime}\,|\alpha\in\Lambda_{i}\}$. Let $R_{\Lambda_{i}}=\cup_{\alpha\in\Lambda_{i}}R_{\alpha}$. If $\mathfrak{Y}=(X,\{R_{\Lambda_{i}}\}_{i=0}^{e})$ becomes an association scheme, it is called a fusion scheme of ${\mathfrak{X}}$, while ${\mathfrak{X}}$ is called a fission scheme of $\mathfrak{Y}$. In particular, $(X,\{R_{0},R_{i}\cup R_{i}^{\top}\}_{i=1}^{d})$ becomes an association scheme, called the symmetrization of $\mathfrak{X}$. If every admissible partition gives rise to a fusion scheme, ${\mathfrak{X}}$ is called amorphic (or amorphous) [14].
In [17], the authors gave the definition of a commutative association scheme generated by a digraph. A digraph $\Gamma$ is a pair $(V(\Gamma),A(\Gamma))$ where $V(\Gamma)$ is a finite nonempty set of vertices and $A(\Gamma)$ is a set of ordered pairs (arcs) $(x,y)$ with distinct vertices $x$ and $y$. For any arc $(x,y)\in A(\Gamma)$, if $A(\Gamma)$ also contains an arc $(y,x)$, then $\{(x,y),(y,x)\}$ can be viewed as an edge. We say that $\Gamma$ is an undirected graph or a graph if $A(\Gamma)$ is a symmetric relation. For an edge $\{(x,y),(y,x)\}$, we say that $x$ is adjacent to $y$, and also call $y$ a neighbour of $x$. A graph is said to be regular of valency $k$ if the number of neighbour of all vertices are equal to $k$. A graph is said to be connected if, for any vertices $x$ and $y$, there is a path from $x$ to $y$. The diameter of a connected graph is the maximum value of the distance function in the graph.
The adjacency matrix $A$ of $\Gamma$ is the $|V(\Gamma)|\times|V(\Gamma)|$ matrix whose $(x,y)$-entry is $1$ if $(x,y)\in A(\Gamma)$, and 00 otherwise. The eigenvalues of a digraph are the eigenvalues of its adjacency matrix. We say that a digraph $\Gamma$ generates a commutative association scheme $\mathfrak{X}$ if and only if the adjacency matrix $A$ of $\Gamma$ generates the Bose-Mesner algebra $\mathcal{M}$ of $\mathfrak{X}$, and in symbols we write $\mathcal{M}=(\langle A\rangle,+,\cdot)$.
In [17], the authors mentioned the following problem.
Problem 1.1.
When can the Bose-Mesner algebra $\mathcal{M}$ of a commutative $d$-class association scheme $\mathfrak{X}$ be generated by a $01$-matrix $A$? In other words, for a given $\mathfrak{X}$, under which combinatorial and algebraic restrictions can we find a $01$-matrix $A$ such that $\mathcal{M}=(\langle A\rangle,+,\cdot)$? Moreover, since such a matrix $A$ is the adjacency matrix of some (di)graph $\Gamma$, can we describe the combinatorial structure of $\Gamma$? The vice-versa question is also of importance, i.e., what combinatorial structure does a (di)graph need to have so that its adjacency matrix will generate the Bose-Mesner algebra of a commutative $d$-class association scheme $\mathfrak{X}$?
In [17], the authors partially answer to questions posted in Problem 1.1. They showed that except for amorphic symmetric association schemes, every $3$-class association scheme can be generated by a (di)graph $\Gamma$ which has $4$ distinct eigenvalues. The authors also described the combinatorial structure of a (di)graph whose adjacency matrix belongs to the Bose-Mesner algebra of a commutative association scheme. At last, the authors gave connections among weakly distance-regular digraphs in the sense of Comellas et al. (see [8] for the details), weakly distance-regular digraphs in the sense of Wang and Suzuki (see [11, 18, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] for the details), and commutative association schemes generated by a $01$-matrix A.
In this paper, we continue to answer the questions in Problem 1.1. The first main theorem is the following result, which give full answers to Problem 1.1 for certain commutative association schemes with exactly one pair of nonsymmetric relations.
Theorem 1.2.
Let $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d+1})$ be a commutative association scheme with $R_{d}^{\top}=R_{d+1}$ and $R_{i}^{\top}=R_{i}$ for $10\leq i\leq d-$ 1, and $\tilde{\mathfrak{X}}$ be its symmetrization. Suppose that the graph $(X,R_{d}\cup R_{d+1})$ generates $\tilde{\mathfrak{X}}$. Then the following hold:
  1. (i)
    The adjacency matrix of $(X,R_{d})$ has exactly $d+2$ distinct eigenvalues;
  2. (ii)
    The digraph $(X,R_{d})$ generates the association scheme $\mathfrak{X}$.
To state our second main theorem, we prepare some more basic notations. Let $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d})$ be a commutative association scheme. Since the Bose-Mesner algebra $\mathcal{M}$ of $\mathfrak{X}$ consists of commuting normal matrices, it has a second basis consisting of primitive idempotents $E_{0}=J/|X|,\dots,E_{d}$. The integers $m_{i}=\textrm{rank}E_{i}$ are called the multiplicities of $\mathfrak{X}$, and $m_{0}=1$ is said to be the trivial multiplicity. For $0\leq i,j\leq d$, there exists a complex scalar $p_{i}(j)$ such that $A_{i}E_{j}=p_{i}(j)E_{j}$ (moreover, $p_{i}(j)$ is the eigenvalue of $A_{i}$ on the eigenspace $V_{j}$), where $E_{j}$ are the orthogonal projectors of $\mathbb{C}^{|X|}$ onto the spaces $V_{j}:=E_{j}\mathbb{C}^{|X|}$. The change-of-basis matrix $P$ is defined by
$A_{i}=\sum_{j=0}^{d}(P)_{ji}E_{j}.$
The $(d+1)\times(d+1)$ matrix $P$ is called the character table or first eigenmatrix of ${\mathfrak{X}}$. According to [2], the character table $P$ has the following form
(5) $\left(\begin{array}[]{cccc}1&k_{1}&\cdots&k_{d}\\ 1&p_{1}(1)&\cdots&p_{d}(1)\\ \vdots&\vdots&\ddots&\vdots\\ 1&p_{1}(d)&\cdots&p_{d}(d)\end{array}\right),$
where the scalars $k_{i},p_{i}(1),\ldots,p_{i}(d)$ are the eigenvalues (not necessarily pairwise distinct) of $A_{i}$ on $V_{0},V_{1},\ldots,V_{d}$, respectively.
The second main theorem concerns a symmetric amorphic association scheme $\tilde{\mathfrak{X}}=(X,\{\tilde{R}_{i}\}_{i=0}^{d})$. Clearly, for each proper subset $\Lambda$ of $\{1,2,\ldots,d\}$, the graph $(X,\tilde{R}_{\Lambda})$ is strongly regular. According to [10, Theorem 1 and Proposition 2], via permutations of the rows and columns, we may assume that the character table $\tilde{\mathfrak{X}}$ has the following form:
(6) $\tilde{P}=\begin{pmatrix}1&k_{1}&k_{2}&k_{3}&\cdots&k_{d}\\ 1&b_{1}&a_{2}&a_{3}&\cdots&a_{d}\\ 1&a_{1}&b_{2}&a_{3}&\cdots&a_{d}\\ 1&a_{1}&a_{2}&b_{3}&\cdots&a_{d}\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&a_{1}&a_{2}&a_{3}&\cdots&b_{d}\end{pmatrix}$
with $a_{i}\neq b_{i}$ for $1\leq i\leq d$.
The second main theorem is as follows, which fully answers Problem 1.1 for commutative association schemes with exactly one pair of nonsymmetric relations whose symmetrizations are amorphic.
Theorem 1.3.
Let $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d+1})$ be a commutative association scheme with $R_{d}^{\top}=R_{d+1}$ and $R_{i}^{\top}=R_{i}$ for $10\leq i\leq d-$ 1, and $\tilde{\mathfrak{X}}$ be its symmetrization with primitive idempotents $\tilde{E}_{0},\tilde{E}_{1},\ldots,\tilde{E}_{d}$. Suppose that $\tilde{\mathfrak{X}}$ is amorphic. Assume that the character table $\tilde{P}$ of $\tilde{\mathfrak{X}}$ has form as (6). Then the following hold:
  1. (i)
    $\mathfrak{X}$ can be generated by a digraph if and only if $d\leq 3$, or $d=4$ and $\tilde{E}_{d}$ is also a primitive idempotent of $\mathfrak{X}$.
  2. (ii)
    If $\mathfrak{X}$ can be generated by a digraph, then there exists $i\in\{1,2,\ldots,d-1,d\}$ such that $(X,R_{i}\cup R_{d})$ has $d+2$ distinct eigenvalues and generates $\mathfrak{X}$.
As an application of Theorems 1.2 and 1.3, we get full answers to questions posted in Problem 1.1 for nonsymmetric $4$-class association schemes, and obtain the third main result as follows.
Theorem 1.4.
Let $\mathfrak{X}=(X,\{R_{0},R_{1},R_{2},R_{3},R_{4}\})$ be a nonsymmetric association scheme with $R_{3}^{\top}=R_{4}$. Then there exists an integer $i\in\{2,3,4\}$ such that the following hold:
  1. (i)
    The adjacency matrix of $(X,R_{i}\cup R_{3})$ has exactly $5$ distinct eigenvalues;
  2. (ii)
    The digraph $(X,R_{i}\cup R_{3})$ generates the association scheme $\mathfrak{X}$.
The remainder of this paper is organized as follows. In Section 2, we recall some basic definitions and results of algebraic graph theory and association schemes. In Section 3, we give proofs of Theorems 1.2 and 1.3. In Section 4, we prove Theorem 1.4 based on Theorems 1.2 and 1.3.
2. Preliminaries
In this section, we shall recall some definitions and results of algebraic graph theory and association schemes which are used frequently in this paper.
2.1. Basic results
First, we recall some results of algebraic graph theory.
Lemma 2.1.
([12, Lemma 8.12.1]) If $\Gamma$ is a connected graph with diameter $d$, then $\Gamma$ has at least $d+1$ distinct eigenvalues.
Lemma 2.2.
([4, Lemma 3.2.1]) Let $\Gamma$ be a regular graph of valency $k$. If $\theta$ is an eigenvalue of $\Gamma$, then $k\geq\theta$.
Proposition 2.3.
([4, Proposition 3.2.2]) Let $\Gamma$ be a regular graph of valency $k$. The multiplicity of the eigenvalue $k$ is the number of connected components of $\Gamma$.
In the remainder of this subsection, we give some results concerning association scheme, and we always assume that $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d})$ is a commutative association scheme with primitive idempotents $E_{0},E_{1}\ldots,E_{d}$ and character table $P$.
Lemma 2.4.
([17, Corollary 2.6]) Let $V_{i}=E_{i}\mathbb{C}^{|X|}$ for $0\leq i\leq d$. Then the sum of the entries of the $V_{i}$ row in $P$ is equal to 00 for $1\leq i\leq d$.
In the rest of this paper, the following theorem, known as the Bannai-Muzychuk criterion for fusion schemes, will be used repeatedly [3, 14].
Theorem 2.5.
Let $\{\Lambda_{i}\}_{i=0}^{e}$ be an admissible partition of the index set $\{0,1,\ldots,d\}$. Then $(X,\{R_{\Lambda_{i}}\}_{i=0}^{e})$ is a fusion scheme if and only if there exists a dual partition $\{\Lambda_{i}^{*}\}_{i=0}^{e}$ of $\{0,1,\ldots,d\}$ with $\Lambda_{0}^{*}=\{0\}$ such that each $(\Lambda_{i}^{*},\Lambda_{j})$ block of the character table $P$ has constant row sum. Moreover, this constant is the $(i,j)$-entry of the character table of this fusion scheme.
We close this subsection with the following result which gives a necessary and sufficient condition for a $01$-matrix generating $\mathfrak{X}$.
Theorem 2.6.
([17, Corollary 2.15]) Let $\mathcal{M}$ be the Bose-Mesner algebra of $\mathfrak{X}$ and $A$ be a $01$-matrix in $\mathcal{M}$. Then $A$ generates $\mathcal{M}$ if and only if $A$ has $d+1$ distinct eigenvalues.
2.2. Symmetric 2-class association schemes: strongly regular graphs
The concept of a strongly regular graph is essentially the same as that of a symmetric $2$-class association scheme. We refer the reader to [4, 5, 6, 12, 19] for further details on the general theory of strongly regular graphs.
Let $(X,R)$ be a regular graph of valency $k$ with $n$ vertices that is neither complete nor empty. Then $(X,R)$ is called strongly regular if any pair of adjacent vertices have $\lambda$ common neighbors, and any two distinct nonadjacent vertices have $\mu$ common neighbors. We say this is a strongly regular graph with parameters $(n,k,\lambda,\mu)$. It is easy to verify that $\mathfrak{X}=(X,\{R_{0},R,\bar{R}\})$ is a symmetric association scheme, where $R_{0}=\{(x,x)\mid x\in X\}$ and $\bar{R}=\{(x,y)\in X\times X\mid(x,y)\notin R\}$.
On the other hand, any symmetric $2$-class association scheme $\mathfrak{X}=(X,\{R_{0},R_{1},R_{2}\})$ is described as a pair of strongly regular graphs $(X,R_{1})$ and $(X,R_{2})$ with parameters $(n,k_{1},p_{1,1}^{1},p_{1,1}^{2})$ and $(n,k_{2},p_{2,2}^{2},p_{2,2}^{1})$, respectively.
Now we consider the strongly regular graphs which are not connected.
Theorem 2.7.
([6, Theorem 3.11]) A strongly regular graph is not connected if and only if it is a disjoint union of at least two complete graphs of the same size.
Proposition 2.8.
Let $(X,R)$ be a strongly regular graph of valency $k$. Then the following are equivalent:
  1. (i)
    $(X,R)$ is not connected;
  2. (ii)
    $(X,R)$ has $-1$ as its eigenvalue;
  3. (iii)
    $(X,R)$ has exactly two eigenvalues: $k$ and $-1$.
Proof.
It is immediate from Theorem 2.7 and [4, Theorem 1.3.1]. ∎
Now let $\Gamma$ be a connected strongly regular graph with parameters $(n,k,\lambda,\mu)$. Counting in two different ways the edges between vertices which are adjacent and nonadjacent to a fixed $x\in X$, we get the well-known identity
(7) $k(k-\lambda-1)=\mu(n-k-1).$
The following result concerns the eigenvalues of a connected strongly regular graph.
Theorem 2.9.
([4, Theorem 1.3.1]) and [12, Lemma 10.2.1] A connected regular graph is strongly regular if and only if it has exactly three distinct eigenvalues.
By Proposition 2.3, Theorem 2.9 and [12, Section 10.2], $\Gamma$ has exactly three eigenvalues: $k$, $r$ and $s$ of multiplicity, respectively, 1, $m_{1}$ and $m_{2}$, where
(8) $r,s$ $=\frac{(\lambda-\mu)\pm\sqrt{(\lambda-\mu)^{2}+4(k-\mu)}}{2},$
(9) $m_{1},m_{2}$ $=\frac{1}{2}\Big{[}(n-1)\mp\frac{2k-(n-1)(\mu-\lambda)}{\sqrt{(\lambda-\mu)^{2}+4(k-\mu)}})\Big{]}.$
By (8), Lemma 2.2 and Proposition 2.8, we have $k>r\geq 0>s\neq-1$. We say that $\Gamma$ is a conference graph if $m_{1}=m_{2}$.
Proposition 2.10.
([12, Lemma 10.3.3]) Let $\Gamma$ be a strongly regular graph with eigenvalues $k$, $r$ and $s$. Then either
  1. (i)
    $\Gamma$ is a conference graph, or
  2. (ii)
    $r$ and $s$ are integers.
On the other hand, we can express the parameters of a connected strongly regular graph $\Gamma$ starting from its eigenvalues and multiplicities: $n=1+m_{1}+m_{2}$,
(10) $\lambda$ $=k+rs+r+s,$
(11) $\mu$ $=k+rs.$
3. Proofs of Theorems 1.2 and 1.3
In this section, we always assume that $\mathfrak{X}=(X,\{R_{i}\}_{i=0}^{d+1})$ is a commutative association scheme with primitive idempotents $E_{0},E_{1},\ldots,E_{d+1}$, where $R_{d}^{\top}=R_{d+1}$ and $R_{i}^{\top}=R_{i}$ for $10\leq i\leq d-$ 1. Denote $\tilde{\mathfrak{X}}=(X,\{\tilde{R}_{i}\}_{i=0}^{d})$ be the symmetrization of $\mathfrak{X}$ with primitive idempotents $\tilde{E}_{0},\tilde{E}_{1},\ldots,\tilde{E}_{d}$, where $\tilde{R}_{d}=R_{d}\cup R_{d+1}$ and $\tilde{R}_{i}=R_{i}$ for $10\leq i\leq d-$ 1.
The following result determines the character table of a commutative association scheme with exactly one pair of nonsymmetric relations.
Theorem 3.1.
([7, Theorem 2.1]) Suppose that the character table $\tilde{P}$ of $\tilde{\mathfrak{X}}$ has form as (5). If $\tilde{E}_{1}=E_{1}+E_{2}$ and $\tilde{E}_{i}=E_{i+1}$ for $2\leq i\leq d$, then the character table of $\mathfrak{X}$ has the following form:
(19) $P=\left(\begin{array}[]{ccccc|cc}1&k_{1}&k_{2}&\cdots&k_{d-1}&k_{d}/2&k_{d}/2\\ \hline\cr 1&p_{1}(1)&p_{2}(1)&\cdots&p_{d-1}(1)&\rho&\bar{\rho}\\ 1&p_{1}(1)&p_{2}(1)&\cdots&p_{d-1}(1)&\bar{\rho}&\rho\\ \hline\cr 1&p_{1}(2)&p_{2}(2)&\cdots&p_{d-1}(2)&p_{d}(2)/2&p_{d}(2)/2\\ 1&p_{2}(3)&p_{3}(3)&\cdots&p_{d-1}(3)&p_{d}(3)/2&p_{d}(3)/2\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 1&p_{2}(d)&p_{3}(d)&\cdots&p_{d-1}(d)&p_{d}(d)/2&p_{d}(d)/2\end{array}\right),$
where $\rho=(p_{d}(1)+\sqrt{a})/2$ and $a<0$.
Now we are ready to give a proof of Theorem 1.2.
Proof of Theorem 1.2.
Suppose that the character table $\tilde{P}$ of $\tilde{\mathfrak{X}}$ has form as (5). By Theorem 2.5, we may assume $\tilde{E}_{1}=E_{1}+E_{2}$ and $\tilde{E}_{j}=E_{j+1}$ for $2\leq j\leq d$. Since $(X,\tilde{R}_{d})$ generates $\tilde{\mathfrak{X}}$, from Theorem 2.6, $(X,\tilde{R}_{d})$ has $d+1$ distinct eigenvalues. In view of (5), all distinct eigenvalues of $(X,\tilde{R}_{d})$ are $k_{d},p_{d}(1),p_{d}(2),\ldots,p_{d}(d)$. By Theorem 3.1, all distinct eigenvalues of $(X,R_{d})$ are $k_{d}/2,p_{d}(2)/2,p_{d}(3)/2,\ldots,p_{d}(d)/2$ and $(p_{d}(1)\pm\sqrt{a})/2$ with $a<0$. Then (i) holds, and (ii) is valid from Theorem 2.6. ∎
In the remainder of this section, we always assume that $\tilde{\mathfrak{X}}$ is an amorphic association scheme with the character table $\tilde{P}$ having form as (6). To give a proof of Theorem 1.3, we need several auxiliary results.
Lemma 3.2.
Let $d\geq 3$. If $1\leq i<j\leq d$, then $a_{i}+b_{j}=a_{j}+b_{i}$.
Proof.
By (6), all eigenvalues of the graph $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ are $k_{i}+k_{j}$, $a_{i}+a_{j}$, $a_{i}+b_{j}$ and $b_{i}+a_{j}$. Since $a_{i}\neq b_{i}$ and $a_{j}\neq b_{j}$, we have $a_{i}+a_{j}\notin\{a_{i}+b_{j},b_{i}+a_{j}\}$.
Suppose, to the contrary that $a_{i}+b_{j}\neq b_{i}+a_{j}$. Then the graph $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ has at least three distinct eigenvalues. Since $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ is a strongly regular graph of valency $k_{i}+k_{j}$, from Proposition 2.8, $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ is connected, which implies that $k_{i}+k_{j}$ is an eigenvalue of $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ with multiplicity $1$ by Proposition 2.3. It follows that $k_{i}+k_{j}\notin\{a_{i}+a_{j},a_{i}+b_{j},b_{i}+a_{j}\}$. Then the connected strongly regular graph $(X,\tilde{R}_{i}\cup\tilde{R}_{j})$ has four distinct eigenvalues, contrary to Theorem 2.9. ∎
Lemma 3.3.
Let $\Lambda$ be a nonempty subset of $\{1,2,\ldots,d+1\}$ with $d\geq 2$. If the digraph $(X,R_{\Lambda})$ has $d+2$ distinct eigenvalues, then exactly one element in $\{d,d+1\}$ belongs to $\Lambda$.
Proof.
Since $(X,\tilde{R}_{\Omega})$ is strongly regular for all proper subset $\Omega$ of $\{1,2,\ldots,d\}$, from Proposition 2.8 and Theorem 2.9, $(X,\tilde{R}_{\Omega})$ has at most three distinct eigenvalues. Thus, the desired result follows. ∎
Lemma 3.4.
Let $V_{i}=E_{i}\mathbb{C}^{|X|}$ for $10\leq i\leq d+$ 1. Assume $E_{i},E_{j}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$ with $11\leq i,j\leq d+$ 1. Suppose that the eigenvalues of $A_{r}$ (resp. $A_{s}$) on the eigenspaces $V_{i}$ and $V_{j}$ are $a_{r}$ and $b_{r}$ (resp. $b_{s}$ and $a_{s}$) respectively with $11\leq r,s\leq d-$ 1. Then $A_{r}$ and $A_{s}$ are the only two adjacency matrices of $\mathfrak{X}$ having distinct eigenvalues on the eigenspaces $V_{i}$ and $V_{j}$.
Proof.
It is immediate from (6) and Theorems 2.5, 3.1. ∎
Lemma 3.5.
Let $V_{i}=E_{i}\mathbb{C}^{|X|}$ for $10\leq i\leq d+$ 1. Suppose that $\mathfrak{X}$ can be generated by a digraph. Then there exist at most two distinct integers $i,j\in\{1,2,\ldots,d+1\}$ such that the eigenvalues of $A_{d}$ on the eigenspaces $V_{i}$ and $V_{j}$ are same.
Proof.
Let $\Gamma$ be a digraph generating $\mathfrak{X}$. Since the adjacency matrix of $\Gamma$ belongs to the Bose-Mesner algebra of $\mathfrak{X}$, there exists a nonempty subset $\Lambda$ of $\{1,2,\ldots,d+1\}$ such that $\Gamma=(X,R_{\Lambda})$. By Theorem 2.6, $(X,R_{\Lambda})$ has $d+2$ distinct eigenvalues.
Assume the contrary, namely, there exist three distinct positive integers $i,j,l\in\{1,2,\ldots,d+1\}$ such that all the eigenvalues of $A_{d}$ on the eigenspaces $V_{i},V_{j}$ and $V_{l}$ are same. By (6) and Theorems 2.5, 3.1, we have $E_{i},E_{j},E_{l}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$, and the eigenvalues of $A_{d}$ on the eigenspaces $V_{i},V_{j}$ and $V_{l}$ are all $a_{d}/2$. Moreover, the eigenvalues of $A_{d+1}$ on the eigenspaces $V_{i},V_{j}$ and $V_{l}$ are also all $a_{d}/2$. By Lemma 3.3, we may assume $d\in\Lambda$. Then $d+1\notin\Lambda$.
Since the digraph $(X,R_{\Lambda})$ has $d+2$ distinct eigenvalues, the eigenvalues of $\sum_{h\in\Lambda}A_{h}$ on the eigenspaces $V_{i}$ and $V_{j}$ are distinct. Since $E_{i},E_{j}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$, from Lemma 3.4, we may assume that $r\in\Lambda$ and the eigenvalues of $A_{r}$ on the eigenspaces $V_{i}$ and $V_{j}$ are $a_{r}$ and $b_{r}$, respectively. Since $E_{l}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$, from (6), the eigenvalue of $A_{r}$ on the eigenspace $V_{l}$ is $a_{r}$. Thus, the eigenvalues of $A_{r}+A_{d}$ on the eigenspaces $V_{i},V_{j}$ and $V_{l}$ are $a_{r}+a_{d}/2$, $b_{r}+a_{d}/2$ and $a_{r}+a_{d}/2$, respectively.
Since the digraph $(X,R_{\Lambda})$ has $d+2$ distinct eigenvalues, the eigenvalues of $\sum_{h\in\Lambda}A_{h}$ on the eigenspaces $V_{i}$ and $V_{l}$ are distinct. Since $E_{i},E_{l}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$, from Lemma 3.4, we may assume that $s\in\Lambda$ and the eigenvalues of $A_{s}$ on the eigenspaces $V_{i}$ and $V_{l}$ are $a_{s}$ and $b_{s}$, respectively. Then $r\neq s$. Since $E_{j}\in\{\tilde{E}_{1},\tilde{E}_{2},\ldots,\tilde{E}_{d-1}\}$, from (6), the eigenvalue of $A_{s}$ on the eigenspace $V_{j}$ is $a_{s}$. Thus, the eigenvalues of $A_{r}+A_{s}+A_{d}$ on the eigenspaces $V_{i},V_{j}$ and $V_{l}$ are $a_{r}+a_{s}+a_{d}/2$, $b_{r}+a_{s}+a_{d}/2$ and $a_{r}+b_{s}+a_{d}/2$, respectively.
By Lemma 3.2, the eigenvalues of $A_{r}+A_{s}+A_{d}$ on the eigenspaces $V_{j}$ and $V_{l}$ are same. Note that the eigenvalues of $A_{r}$ (resp. $A_{s}$) on the eigenspaces $V_{j}$ and $V_{l}$ are $b_{r}$ and $a_{r}$ (resp. $a_{s}$ and $b_{s}$), respectively. By Lemma 3.4, the eigenvalues of $A_{t}$ on the eigenspaces $V_{j}$ and $V_{l}$ are same for all $t\in\Lambda\setminus\{s,r\}$. It follows that the eigenvalues of $\sum_{h\in\Lambda}A_{h}$ on the eigenspaces $V_{j}$ and $V_{l}$ are also same, contrary to the fact that $(X,R_{\Lambda})$ has $d+2$ distinct eigenvalues. ∎
Lemma 3.6.
If $d\geq 3$, then $a_{i},b_{i}<k_{i}$ for $1\leq i\leq d$.
Proof.
By [10, Theorem 1], $(X,\tilde{R}_{i})$ is a connected strongly regular graph for $1\leq i\leq d$. By (6), Lemma 2.2 and Theorem 2.9, we have $a_{i},b_{i}<k_{i}$. ∎
Proposition 3.7.
Let $d=3$. If $\tilde{E}_{3}\notin\{E_{1},E_{2},E_{3},E_{4}\}$, then $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues for $i\in\{1,2\}$.
Proof.
By Theorem 2.5, we may assume $\tilde{E}_{3}=E_{3}+E_{4}$ and $\tilde{E}_{h}=E_{h}$ for $0\leq h\leq 2$. In view of (6) and Theorem 3.1, the character table of $\mathfrak{X}$ has the form
(25) $P=\left(\begin{array}[]{ccccc}1&k_{1}&k_{2}&k_{3}/2&k_{3}/2\\ 1&b_{1}&a_{2}&a_{3}/2&a_{3}/2\\ 1&a_{1}&b_{2}&a_{3}/2&a_{3}/2\\ 1&a_{1}&a_{2}&(b_{3}+\sqrt{a})/2&(b_{3}-\sqrt{a})/2\\ 1&a_{1}&a_{2}&(b_{3}-\sqrt{a})/2&(b_{3}+\sqrt{a})/2\end{array}\right)$
with $a<0$. Then eigenvalues of $(X,R_{i}\cup R_{3})$ are $k_{i}+k_{3}/2$, $a_{i}+a_{3}/2$, $b_{i}+a_{3}/2$ and $(2a_{i}+b_{3}\pm\sqrt{a})/2$. Since $a_{i}\neq b_{i}$, from Lemma 3.6, $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues. ∎
Proposition 3.8.
Suppose $\tilde{E}_{d}\in\{E_{1},E_{2},\ldots,E_{d+1}\}$. The following hold:
  1. (i)
    If $d=3$, then $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues for all $i\in\{1,2,3\}$;
  2. (ii)
    If $d=4$, then $(X,R_{i}\cup R_{4})$ has $6$ distinct eigenvalues for all $i\in\{2,3\}$.
Proof.
In view of (6) and Theorem 2.5, we may assume $\tilde{E}_{1}=E_{1}+E_{2}$ and $\tilde{E}_{h}=E_{h+1}$ for $2\leq h\leq d$. By Theorem 3.1, the character table of $\mathfrak{X}$ has the form
(33) $P=\left(\begin{array}[]{ccccccc}1&k_{1}&k_{2}&\cdots&k_{d-1}&k_{d}/2&k_{d}/2\\ 1&b_{1}&a_{2}&\cdots&a_{d-1}&(a_{d}+\sqrt{a})/2&(a_{d}-\sqrt{a})/2\\ 1&b_{1}&a_{2}&\cdots&a_{d-1}&(a_{d}-\sqrt{a})/2&(a_{d}+\sqrt{a})/2\\ 1&a_{1}&b_{2}&\cdots&a_{d-1}&a_{d}/2&a_{d}/2\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 1&a_{1}&a_{2}&\cdots&b_{d-1}&a_{d}/2&a_{d}/2\\ 1&a_{1}&a_{2}&\cdots&a_{d-1}&b_{d}/2&b_{d}/2\end{array}\right)$
with $a<0$.
(i) Since $a_{3}\neq b_{3}$, from (33) and Lemma 3.6, all distinct eigenvalues of $(X,R_{3})$ are $k_{3}/2$, $(a_{3}\pm\sqrt{a})/2$, $a_{3}/2$ and $b_{3}/2$.
By (33), $(X,R_{1}\cup R_{3})$ has eigenvalues $k_{1}+k_{3}/2$, $(2b_{1}+a_{3}\pm\sqrt{a})/2$, $a_{1}+a_{3}/2$ and $a_{1}+b_{3}/2$. Since $a_{3}\neq b_{3}$, from Lemma 3.6, $(X,R_{1}\cup R_{3})$ has $5$ distinct eigenvalues.
By (33), $(X,R_{2}\cup R_{3})$ has eigenvalues $k_{2}+k_{3}/2$, $(2a_{2}+a_{3}\pm\sqrt{a})/2$, $b_{2}+a_{3}/2$ and $a_{2}+b_{3}/2$. Since $a_{3}\neq b_{3}$ and $b_{2}+a_{3}=a_{2}+b_{3}$ from Lemma 3.2, we have $b_{2}+a_{3}/2\neq a_{2}+b_{3}/2$. By Lemma 3.6, $(X,R_{2}\cup R_{3})$ has $5$ distinct eigenvalues. Thus, (i) is valid.
(ii) By (33), $(X,R_{i}\cup R_{4})$ has eigenvalues $k_{i}+k_{4}/2$, $a_{i}+a_{4}/2$, $b_{i}+a_{4}/2$, $a_{i}+b_{4}/2$ and $(2a_{i}+a_{4}\pm\sqrt{a})/2$. Since $a_{h}\neq b_{h}$ for $1\leq h\leq d$ and $b_{i}+a_{4}=a_{i}+b_{4}$ from Lemma 3.2, $a_{i}+a_{4}/2$, $b_{i}+a_{4}/2$ and $a_{i}+b_{4}/2$ are distinct. By Lemma 3.6, $(X,R_{i}\cup R_{4})$ has $6$ distinct eigenvalues. Thus, (ii) is also valid. ∎
Now we are ready to prove Theorem 1.3.
Proof of Theorem 1.3.
If $d=1$, from Theorems 2.5 and 3.1, then the graph $(X,R_{j})$ has exactly $3$ distinct eigenvalues for $j\in\{1,2\}$, which implies that $(X,R_{j})$ generates $\mathfrak{X}$ by Theorem 2.6. Then the sufficiency of (i), and (ii) are both valid from [17, Theorem 3.8], Propositions 3.7, 3.8 and Theorem 2.6.
It suffices to prove the necessity of (i). Let $\Gamma$ be a digraph generating $\mathfrak{X}$. Since the adjacency matrix of $\Gamma$ belongs to the Bose-Mesner algebra of $\mathfrak{X}$, there exists a proper subset $\Lambda$ of $\{1,2,\ldots,d+1\}$ such that $\Gamma=(X,R_{\Lambda})$. Assume the contrary, namely, $d\geq 5$, or $d=4$ and $\tilde{E}_{4}\notin\{E_{1},E_{2},E_{3},E_{4}\}$. Let $V_{h}=E_{h}\mathbb{C}^{|X|}$ for $10\leq h\leq d+$ 1.
Suppose $\tilde{E}_{d}\notin\{E_{1},E_{2},\ldots,E_{d+1}\}$. By Theorem 2.5, without loss of generality, we may assume $\tilde{E}_{d}=E_{d}+E_{d+1}$ and $\tilde{E}_{h}=E_{h}$ for $10\leq i\leq d-$ 1. Since $d\geq 4$, from (6) and Theorem 3.1, the eigenvalues of $A_{d}$ on the eigenspaces $V_{1}$, $V_{2}$ and $V_{3}$ are all $a_{d}/2$, contrary to Lemma 3.5. Then, $d\geq 5$ and $\tilde{E}_{d}\in\{E_{1},E_{2},\ldots,E_{d+1}\}$.
By Theorem 2.5, we may assume $\tilde{E}_{1}=E_{1}+E_{2}$ and $\tilde{E}_{h}=E_{h+1}$ for $2\leq i\leq d$. Since $d\geq 5$, from (6) and Theorem 3.1, the eigenvalues of $A_{d}$ on the eigenspaces $V_{3}$, $V_{4}$ and $V_{5}$ are all $a_{d}/2$, contrary to Lemma 3.5. Thus, the necessity of (i) is also valid. ∎
4. Proof of Theorem 1.4
In this section, we always assume that $\mathfrak{X}=(X,\{R_{0},R_{1},R_{2},R_{3},R_{4}\})$ is an association scheme with $R_{3}^{\top}=R_{4}$. By a theorem of Higman [13], association schemes with at most four classes are commutative. Next, we divide this section into two subsections according to whether $\mathfrak{X}$ is skew-symmetric.
4.1. $\mathfrak{X}$ is skew-symmetric
Since $\mathfrak{X}$ is skew-symmetric, we have $R_{1}^{\top}=R_{2}$. At first, we recall the character table of $\mathfrak{X}$.
Theorem 4.1.
([15, Theorem 5]) Let $\tilde{\mathfrak{X}}=(X,\{R_{0},R_{1}\cup R_{2},R_{3}\cup R_{4}\})$ be the symmetrization of $\mathfrak{X}$, and
(40) $\tilde{P}=\left(\begin{array}[]{ccc}1&k_{1}&k_{2}\\ 1&r_{1}&t_{1}\\ 1&r_{2}&t_{2}\end{array}\right)\begin{array}[]{c}1\\ m_{1}\\ m_{2}\end{array}$
be the character table of $\tilde{\mathfrak{X}}$, where $1,m_{1}$ and $m_{2}$ are the multiplicities of the corresponding row eigenvalues. Then the character table of $\mathfrak{X}$ has the following form::::
$P=\left(\begin{array}[]{ccccc}1&k_{1}/2&k_{1}/2&k_{2}/2&k_{2}/2\\ 1&\rho&\bar{\rho}&\tau&\bar{\tau}\\ 1&\sigma&\bar{\sigma}&\omega&\bar{\omega}\\ 1&\bar{\sigma}&\sigma&\bar{\omega}&\omega\\ 1&\bar{\rho}&\rho&\bar{\tau}&\tau\end{array}\right)\begin{array}[]{c}1\\ m_{1}/2\\ m_{1}/2\\ m_{2}/2\\ m_{2}/2\end{array}.$
Let $n=k_{1}+k_{2}+1$. The entries $\rho,\omega,\tau,$ and $\sigma$ are one of the three cases:
  1. (i)
    $\rho=r_{1}/2,\ \sigma=(r_{2}+\sqrt{-b})/2,\ \tau=(t_{1}+\sqrt{-z})/2,\ \omega=t_{2}/2,$ where $b={nk_{1}}/{m_{2}},z={nk_{2}}/{m_{1}}.$
  2. (ii)
    $\rho=\left(r_{1}+\sqrt{-y}\right)/2,\ \sigma={r_{2}}/{2},\ \tau={t_{1}}/{2},\ \omega=\left(t_{2}+\sqrt{-c}\right)/2,$ where $y={nk_{1}}/{m_{1}},c={nk_{2}}/{m_{2}}.$
  3. (iii)
    $\rho=(r_{1}+\sqrt{-y})/2,\ \tau=(t_{1}+\sqrt{-z})/2,\ \sigma=(r_{2}+\sqrt{-b})/2,\ \omega=(t_{2}-\sqrt{-c})/2,$ where all of $b,c,y,z$ are positive.
In this subsection, we refer to the three character tables in the above theorem as type 1, 2, and 3, respectively.
Proposition 4.2.
If $\mathfrak{X}$ has the character table of type $h$ for some $h\in\{1,2\}$, then the digraph $(X,R_{i}\cup R_{i+2})$ has $5$ distinct eigenvalues for $i\in\{1,2\}$.
Proof.
Let $\tilde{\mathfrak{X}}=(X,\{R_{0},R_{1}\cup R_{2},R_{3}\cup R_{4}\})$ be the symmetrization of $\mathfrak{X}$ having the character table $\tilde{P}$ as (40). Since both the graphs $(X,R_{1}\cup R_{2})$ and $(X,R_{3}\cup R_{4})$ are strongly regular, from Theorem 2.7, we may assume that $(X,R_{1}\cup R_{2})$ is connected. Let $(X,R_{1}\cup R_{2})$ have parameters $(n,k,\lambda,\mu)$ with $k=k_{1}$. Since $(X,R_{1}\cup R_{2})$ has eigenvalues $k,r_{1},r_{2}$ from (40), one has $k=\mu-r_{1}r_{2}$ by (11). In view of (10), we get $\lambda=\mu+r_{1}+r_{2}$. Since $(X,R_{1}\cup R_{2})$ is connected, one obtains $\mu>0$. By substituting $k=\mu-r_{1}r_{2}$ and $\lambda=\mu+r_{1}+r_{2}$ into (7), we have
(41) $n=\frac{(\mu-r_{1}r_{2}-r_{2})(\mu-r_{1}r_{2}-r_{1})}{\mu}.$
This implies
(42) $k_{2}=\frac{(r_{2}+1)(r_{1}+1)(r_{1}r_{2}-\mu)}{\mu}.$
By substituting (41), $k=\mu-r_{1}r_{2}$ and $\lambda=\mu+r_{1}+r_{2}$ into (9), one gets
(43) $m_{1}$ $=\frac{(r_{2}+1)(\mu-r_{1}r_{2}-r_{2})(\mu-r_{1}r_{2})}{\mu(r_{2}-r_{1})},$
(44) $m_{2}$ $=\frac{(r_{1}+1)(\mu-r_{1}r_{2}-r_{1})(\mu-r_{1}r_{2})}{\mu(r_{1}-r_{2})}.$
By Theorem 4.1 (i) and (ii), the digraph $(X,R_{i}\cup R_{i+2})$ has eigenvalues $(k_{1}+k_{2})/2,\rho+\tau,\bar{\rho}+\bar{\tau},\sigma+\omega$ and $\bar{\sigma}+\bar{\omega}$ for $i\in\{1,2\}$, where $\rho+\tau=(r_{1}+t_{1}+\sqrt{-nk_{l}/m_{1}})/2$ and $\sigma+\omega=(r_{2}+t_{2}+\sqrt{-nk_{h}/m_{2}})/2$ with $\{l,h\}=\{1,2\}$. Since all of $n,k_{1},k_{2},m_{1},m_{2}$ are positive, from Theorem 2.6, it suffices to show that $\rho+\tau\neq\sigma+\omega$.
Assume the contrary, namely, $\rho+\tau=\sigma+\omega$. It follows that
$(r_{1}+t_{1}+\sqrt{-nk_{l}/m_{1}})/2=\rho+\tau=\sigma+\tau=(r_{2}+t_{2}+\sqrt{-nk_{h}/m_{2}})/2,$
and so $k_{l}m_{2}=k_{h}m_{1}$. By substituting $k=\mu-r_{1}r_{2}$ and (42)–(44) into the above equation, one gets
(45) $\frac{(\mu-r_{1}r_{2})^{2}(r_{l}+1)(r_{h}^{2}+r_{h}-\mu)(\mu-r_{1}r_{2}-r_{1}-r_{2}-1)}{\mu^{2}(r_{1}-r_{2})}=0.$
By Subsection 2.2 and Proposition 2.8, we obtain $r_{1}r_{2}\leq 0$ and $-1\notin\{r_{1},r_{2}\}$. Since $k_{2}>0$, from (42), one has $(r_{2}+1)(r_{1}+1)<0$. (45) implies $\mu=r_{h}^{2}+r_{h}$. Since $\lambda=\mu+r_{1}+r_{2}$, we get $\lambda=r_{h}^{2}+2r_{h}+r_{l}$.
In view of [16, Theorem 2.2 (i) and (ii)], we get $p_{1,3}^{1}-p_{1,3}^{2}=t_{l}/2$ and $p_{1,3}^{3}-p_{1,3}^{4}=r_{h}/2$, which imply that both $t_{l}$ and $r_{h}$ are even. Since $01+r_{l}+t_{l}=$ 0 from Lemma 2.4, $r_{l}$ is odd. Since $\lambda=r_{h}^{2}+2r_{h}+r_{l}$, from [16, Theorem 2.2 (i) and (ii)] again, we obtain
$p_{1,1}^{1}=\frac{\lambda+r_{l}}{4}=\frac{r_{h}(r_{h}+2)+2r_{l}}{4}.$
The fact that $r_{h}$ is even implies $2\mid r_{l}$, a contradiction. ∎
Proposition 4.3.
If $\mathfrak{X}$ has the character table of type 3, then the digraph $(X,R_{i})$ has $5$ distinct eigenvalues for $1\leq i\leq 4$.
Proof.
Let the symmetrization of $\mathfrak{X}$ have the character table $\tilde{P}$ as (40). By Theorem 4.1 (iii), the digraph $(X,R_{i})$ with $i\in\{1,2\}$ has eigenvalues $k_{1}/2,(r_{1}\pm\sqrt{-y})/2$ and $(r_{2}\pm\sqrt{-b})/2$, and the digraph $(X,R_{j})$ with $j\in\{3,4\}$ has eigenvalues $k_{2}/2,(t_{1}\pm\sqrt{-z})/2$ and $,(t_{2}\pm\sqrt{-c})/2$, where all of $b,c,y,z$ are positive. By Theorem 2.6, it suffices to show that $r_{1}\neq r_{2}$ and $t_{1}\neq t_{2}$.
Since $\tilde{P}$ is a change-of-basis matrix between the adjacency matrices of $\tilde{\mathfrak{X}}$ and the primitive idempotents of its Bose-Mesner algebra, $\tilde{P}$ is not singular. By Lemma 2.4 and (40), one has $01+r_{1}+t_{1}=1+r_{2}+t_{2}=$ 0, which implies $r_{1}\neq r_{2}$ and $t_{1}\neq t_{2}$. ∎
4.2. $\mathfrak{X}$ is not skew-symmetric
Since $\mathfrak{X}$ is not skew-symmetric, we have $R_{1}^{\top}=R_{1}$ and $R_{2}^{\top}=R_{2}$. Let $\tilde{\mathfrak{X}}=(X,\{\tilde{R}_{i}\}_{i=0}^{3})$ be its symmetrization with $\tilde{R}_{3}=R_{3}\cup R_{4}$ and $\tilde{R}_{i}=R_{i}$ for $0\leq i\leq 2$. Set $\Gamma^{(i)}=(X,\tilde{R}_{i})$ for $1\leq i\leq 3$. In [9], van Dam completely classified symmetric $3$-class association schemes.
Lemma 4.4.
([9, Section 7]) Suppose that $\tilde{\mathfrak{X}}$ is not amorphic. Then $\tilde{\mathfrak{X}}$ belongs to exactly one of the following categories:
  1. (i)
    $3$-class association schemes with at least one graph $\Gamma^{(i)}$ that is the disjoint union of $N>1$ ($N\in\mathbb{N}$) connected strongly regular graphs with the same parameters;
  2. (ii)
    $3$-class association schemes with at least one graph $\Gamma^{(i)}$ having $4$ distinct eigenvalues.
Let $E_{0},E_{1},E_{2},E_{3},E_{4}$ be primitive idempotents of $\mathfrak{X}$, and $\tilde{E}_{0},\tilde{E}_{1},\tilde{E}_{2},\tilde{E}_{3}$ be primitive idempotents of $\tilde{\mathfrak{X}}$. In Lemma 4.5, the shape of the character table of the first class of graphs from Lemma 4.4 is described.
Lemma 4.5.
([17, Lemma 3.3]) Suppose that $\Gamma^{(1)}$ is the disjoint union of $N>1$ $(N\in\mathbb{N}$) connected strongly regular graphs, each one with valency $k$, eigenvalues $r$ and $s$ $(r\neq k,s\neq k)$, and $w$ vertices. Then the character table $\tilde{\mathfrak{X}}$ has the form
(50) $\tilde{P}=\left(\begin{array}[]{cccc}1&k&w-k-1&(N-1)w\\ 1&k&w-k-1&-w\\ 1&r&-1-r&0\\ 1&s&-1-s&0\end{array}\right).$
Proposition 4.6.
Suppose that $\Gamma^{(i)}$ for some $i\in\{1,2\}$ is the disjoint union of $N>1$ connected strongly regular graphs, each one with valency $k$, eigenvalues $r$ and $s$ $(r\neq k,s\neq k)$, and $w$ vertices. Then $(X,R_{3})$ or $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues.
Proof.
Without loss of generality, we may assume $i=1$. In view of Lemma 2.2 and Theorem 2.9, we have $k>r,s$ and $r\neq s$. By Lemma 4.5, we may assume that the character table $\tilde{P}$ of $\tilde{\mathfrak{X}}$ has the form as (50). By replacing the roles of $\tilde{E}_{2}$ and $\tilde{E}_{3}$, from Theorem 2.5, we may assume that $\tilde{E}_{1}=E_{1}+E_{2}$ and $\tilde{E}_{l}=E_{l+1}$ for $l\in\{2,3\}$, or $\tilde{E}_{3}=E_{3}+E_{4}$ and $\tilde{E}_{l}=E_{l}$ for $l\in\{1,2\}$.
Suppose $\tilde{E_{1}}=E_{1}+E_{2}$ and $\tilde{E}_{l}=E_{l+1}$ for $l\in\{2,3\}$. By (50) and Theorem 3.1, the character table of $\mathfrak{X}$ has the following form:
(56) $P=\left(\begin{array}[]{ccccc}1&k&w-k-1&(N-1)w/2&(N-1)w/2\\ 1&k&w-k-1&(-w+\sqrt{a})/2&(-w-\sqrt{a})/2\\ 1&k&w-k-1&(-w-\sqrt{a})/2&(-w+\sqrt{a})/2\\ 1&r&-1-r&0&0\\ 1&s&-1-s&0&0\end{array}\right)$
with $a<0$. Then the digraph $(X,R_{1}\cup R_{3})$ has eigenvalues $k+(N-1)w/2,(2k-w\pm\sqrt{a})/2,r$ and $s$. Since $k>r,s$ and $r\neq s$, the digraph $(X,R_{1}\cup R_{3})$ has $5$ distinct eigenvalues.
Suppose $\tilde{E}_{3}=E_{3}+E_{4}$ and $\tilde{E}_{h}=E_{h}$ for $h\in\{1,2\}$. By (50) and Theorem 3.1, the character table of $\mathfrak{X}$ has the following form:
(62) $P=\left(\begin{array}[]{ccccc}1&k&w-k-1&(N-1)w/2&(N-1)w/2\\ 1&k&w-k-1&-w/2&-w/2\\ 1&r&-1-r&0&0\\ 1&s&-1-s&\sqrt{a}&-\sqrt{a}\\ 1&s&-1-s&-\sqrt{a}&\sqrt{a}\end{array}\right)$
with $a<0$. Then the digraph $(X,R_{3})$ has eigenvalues $(N-1)w/2,\pm\sqrt{a},-w/2$ and 00. Since $N,w>1$, the digraph $(X,R_{3})$ has $5$ distinct eigenvalues. ∎
Proposition 4.7.
Suppose that $\Gamma^{(3)}$ is the disjoint union of $N>1$ connected strongly regular graphs, each one with valency $k$, eigenvalues $r$ and $s$ $(r\neq k,s\neq k)$, and $w$ vertices. Then $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues for some $i\in\{1,2,3\}$.
Proof.
Since $(X,\tilde{R}_{3})$ is a connected strongly regular graph, from Theorem 2.9, $k$ and $r,s$ are distinct. By Lemma 4.5, there exists $i\in\{1,2\}$ such that the valency of $\Gamma^{(i)}$ is not $w-k-1$. Without loss of generality, we may assume $i=2$. By Lemma 4.5, the character table of $\tilde{\mathfrak{X}}$ has the form
(67) $\tilde{P}=\left(\begin{array}[]{cccc}1&w-k-1&(N-1)w&k\\ 1&w-k-1&-w&k\\ 1&-1-r&0&r\\ 1&-1-s&0&s\end{array}\right).$
By replacing the roles of $\tilde{E}_{2}$ and $\tilde{E}_{3}$, from Theorem 2.5, we may assume that $\tilde{E_{1}}=E_{1}+E_{2}$ and $\tilde{E}_{h}=E_{h+1}$ for $h\in\{2,3\}$, or $\tilde{E}_{3}=E_{3}+E_{4}$ and $\tilde{E}_{h}=E_{h}$ for $h\in\{1,2\}$.
Case 1. $\tilde{E_{1}}=E_{1}+E_{2}$ and $\tilde{E}_{h}=E_{h+1}$ for $h\in\{2,3\}$.
By (67) and Theorem 3.1, the character table of $\mathfrak{X}$ has the following form:
(73) $P=\left(\begin{array}[]{ccccc}1&w-k-1&(N-1)w&k/2&k/2\\ 1&w-k-1&-w&(k+\sqrt{a})/2&(k-\sqrt{a})/2\\ 1&w-k-1&-w&(k-\sqrt{a})/2&(k+\sqrt{a})/2\\ 1&-1-r&0&r/2&r/2\\ 1&-1-s&0&s/2&s/2\end{array}\right)$
with $a<0$. Then the digraph $(X,R_{3})$ has eigenvalues $k/2,(k\pm\sqrt{a})/2,r/2$ and $s/2$. Since $k,r$ and $s$ are distinct, the digraph $(X,R_{3})$ has $5$ distinct eigenvalues.
Case 2. $\tilde{E}_{3}=E_{3}+E_{4}$ and $\tilde{E}_{h}=E_{h}$ for $h\in\{1,2\}$.
By (67) and Theorem 3.1, the character table of $\mathfrak{X}$ has the following form:
(79) $P=\left(\begin{array}[]{ccccc}1&w-k-1&(N-1)w&k/2&k/2\\ 1&w-k-1&-w&k/2&k/2\\ 1&-1-r&0&r/2&r/2\\ 1&-1-s&0&(s+\sqrt{a})/2&(s-\sqrt{a})/2\\ 1&-1-s&0&(s-\sqrt{a})/2&(s+\sqrt{a})/2\end{array}\right)$
with $a<0$. Then the digraph $(X,R_{2}\cup R_{3})$ has eigenvalues $k/2+(N-1)w,(s\pm\sqrt{a})/2,k/2-w$ and $r/2$. To prove that $(X,R_{2}\cup R_{3})$ has $5$ distinct eigenvalues, it suffices to show that $w\neq(k-r)/2$ since $r<k$ from Lemma 2.2.
Assume the contrary, namely, $w=(k-r)/2$. Then $r<0$. Since $(X,\tilde{R}_{3})$ is a connected strongly regular graph on $w$ vertices with the parameters $(w,k,\lambda,\mu)$, one obtains $\mu>0$. By (10) and (11), we have $\lambda=\mu+r+s$ and $k=\mu-rs$. In view of (7), one gets
(80) $w=\frac{(\mu-rs-s)(\mu-rs-r)}{\mu}.$
By substituting $w=(k-r)/2=(\mu-rs-r)/2$ into (80), we obtain
$\frac{(\mu-rs-r)(\mu-2rs-2s)}{\mu}=0.$
Since $w>0$, from (80), one has $\mu=2s(r+1)$. By Subsection 2.2, we have $k>s\geq 0>r$. The fact $\mu>0$ implies $-1<r<0<s$. By Proposition 2.10, $(X,\tilde{R}_{3})$ is a conference graph. (9) implies $2k=(w-1)(\mu-\lambda)$. By substituting $w,k,\lambda,\mu$ into the above equation, we get $(r+2)(rs+s^{2}-r+3s)=0$. Since $-1<r<0<s$, we have $r=-\frac{s(s+3)}{s-1}$, and so $s>1$. Then $\mu=2s(r+1)=-2s(s+1)^{2}/(s-1)<0,$ a contradiction. ∎
Proposition 4.8.
Suppose that $\Gamma^{(i)}$ has $4$ distinct eigenvalues for some $i\in\{1,2,3\}$. Then $(X,R_{j}\cup R_{3})$ has $5$ distinct eigenvalues for some $j\in\{1,2,3\}$.
Proof.
By Theorems 1.2 and 2.6, we only need to consider the case that $\Gamma^{(3)}$ does not have $4$ distinct eigenvalues. Without loss of generality, we may assume that $\Gamma^{(1)}$ has $4$ distinct eigenvalues. It follows from Lemma 2.4 that the character table of $\tilde{\mathfrak{X}}$ has the following form:
(85) $\tilde{P}=\left(\begin{array}[]{cccc}1&k_{1}&k_{2}&2k_{3}\\ 1&u&-1-r-u&r\\ 1&v&-1-s-v&s\\ 1&w&-1-t-w&t\end{array}\right),$
where $k_{1},u,v,w$ are distinct and |$|\{2k_{3},r,s,t\}|<4$. Since $(X,\tilde{R}_{1})$ (resp. $(X,\tilde{R}_{3})$) is a graph of valency $k_{1}$ with eigenvalues $k_{1},u,v,w$ (resp. valency $2k_{3}$ with eigenvalues $2k_{3},r,s,t$) from (85), we have $k_{1}>u,v,w$ and $2k_{3}\geq r,s,t$ by Lemma 2.2.
By Theorem 2.5, we may assume $\tilde{E}_{1}=E_{1}+E_{2}$, $\tilde{E}_{2}=E_{3}$ and $\tilde{E}_{3}=E_{4}$. In view of (85) and Theorem 3.1, the character table of $\tilde{X}$ has the following form:
(91) $P=\left(\begin{array}[]{ccccc}1&k_{1}&k_{2}&k_{3}&k_{3}\\ 1&u&-1-r-u&(r+\sqrt{a})/2&(r-\sqrt{a})/2\\ 1&u&-1-r-u&(r-\sqrt{a})/2&(r+\sqrt{a})/2\\ 1&v&-1-s-v&s/2&s/2\\ 1&w&-1-t-w&t/2&t/2\end{array}\right)$
with $a<0$.
Assume the contrary, namely, none of the digraphs $(X,R_{3})$, $(X,R_{1}\cup R_{3})$ and $(X,R_{2}\cup R_{3})$ have $5$ distinct eigenvalues. Suppose $s=t$. By (91), $(X,R_{1}\cup R_{3})$ has eigenvalues $k_{1}+k_{3},(2u+r\pm\sqrt{a})/2,v+s/2$ and $w+s/2$. The fact $w\neq v$ implies $k_{1}+k_{3}\in\{v+s/2,w+s/2\}$, contrary to the fact that $k_{1}>v,w$ and $2k_{3}\geq s$. Then $s\neq t$.
By (91), $(X,R_{3})$ has eigenvalues $k_{3},(r\pm\sqrt{a})/2,s/2$ and $t/2$. Since $(X,R_{3})$ has less than $5$ distinct eigenvalues, we have $2k_{3}\in\{t,s\}$. Without loss of generality, we may assume $s=2k_{3}$. Since all eigenvalues of $\Gamma^{(3)}$ are $2k_{3},r$ and $t$ from (85), by Lemma 2.1 and Theorem 2.9, $\Gamma^{(3)}$ is the disjoint union of complete graphs with the same size, or the disjoint union of connected strongly regular graphs with the same parameters. By Proposition 4.7, $(X,\tilde{R}_{3})$ is the disjoint union of $N>1$ complete graphs $K_{2k_{3}+1}$. Since $t\neq 2k_{3}$, from Theorem 2.9, we have $r\in\{2k_{3},t\}$. In view of Theorem 2.7, the strongly regular graph $\Gamma^{(3)}$ is not connected. Proposition 2.8 implies $t=-1$.
To distinguish the notations between $\mathfrak{X}$ and $\tilde{\mathfrak{X}}$, let $p_{i^{\prime},j^{\prime}}^{h^{\prime}}$ denote the intersection numbers belonging to $\mathfrak{X}$ with $0\leq i^{\prime},j^{\prime},h^{\prime}\leq 4$, and $\tilde{p}_{i^{\prime\prime},j^{\prime\prime}}^{h^{\prime\prime}},\tilde{m}_{i^{\prime\prime}}$ the parameters belonging to $\tilde{\mathfrak{X}}$ with $0\leq i^{\prime\prime},j^{\prime\prime},h^{\prime\prime}\leq 3$. By [7, Theorem 2.6 (iii)], one gets $p_{3,1}^{3}=(\tilde{p}_{3,1}^{3}+u)/2$ and $p_{3,2}^{3}=(\tilde{p}_{3,2}^{3}-1-r-u)/2$. Since $(X,\tilde{R}_{3})$ is the disjoint union of $N>1$ complete graphs $K_{2k_{3}+1}$, we have $p_{3,l}^{3}=\tilde{p}_{3,l}^{3}=0$ for $l\in\{1,2\}$. It follows that $u=0$ and $r=-1$.
By (91), $(X,R_{1}\cup R_{3})$ has eigenvalues $k_{1}+k_{3},(-1\pm\sqrt{a})/2,v+k_{3}$ and $w-1/2$. Since $v,w<k_{1}$ and $(X,R_{1}\cup R_{3})$ has less than $5$ distinct eigenvalues, one has $w=v+k_{3}+1/2$, which implies
(96) $\tilde{P}=\left(\begin{array}[]{cccc}1&k_{1}&k_{2}&2k_{3}\\ 1&0&0&-1\\ 1&v&-1-v-2k_{3}&2k_{3}\\ 1&1/2+v+k_{3}&-1/2-v-k_{3}&-1\end{array}\right).$
Since $\Gamma^{(3)}$ is the disjoint union of $N>1$ complete graphs $K_{2k_{3}+1}$, the configuration $\hat{\mathfrak{X}}:=(X,\{\hat{R}_{0},\hat{R}_{1},\hat{R}_{2}\})$ is a $2$-class symmetric association scheme with $\hat{R}_{1}=\tilde{R}_{1}\cup\tilde{R}_{2}$ and $\hat{R}_{2}=\tilde{R}_{3}$. Let $\hat{E}_{0},\hat{E}_{1},\hat{E}_{2}$ be the primitive idempotents and $\hat{m}_{0},\hat{m}_{1},\hat{m}_{2}$ be multiplicities of $\hat{\mathfrak{X}}$. Since the strongly regular graph $\Gamma^{(3)}$ is not connected, from Lemma 2.4 and Propositions 2.3, 2.8, the character table $\hat{\mathfrak{X}}$ has the form:
(103) $\hat{P}=\left(\begin{array}[]{cccc}1&(N-1)(2k_{3}+1)&2k_{3}\\ 1&0&-1\\ 1&-1-2k_{3}&2k_{3}\end{array}\right)\begin{array}[]{c}1\\ \hat{m}_{1}=2Nk_{3}\\ \hat{m}_{2}=N-1\end{array},$
where $k_{1}+k_{2}=(N-1)(2k_{3}+1)$. By Theorem 2.5 and (96), one has $\tilde{E}_{1}+\tilde{E}_{3}=\hat{E}_{1}$ and $\tilde{E}_{2}=\hat{E}_{2}$, which imply $\tilde{m}_{1}+\tilde{m}_{3}=\hat{m}_{1}=2Nk_{3}$ and $\tilde{m}_{2}=\hat{m}_{2}=N-1$. In view of [3, (2.1)], we get $v=-(2k_{3}+1)k_{1}/(k_{1}+k_{2})=-k_{1}/(N-1)$. Since the trace of the adjacency matrix of $\tilde{R}_{1}$ is 00, one obtains
$k_{1}+v\tilde{m}_{2}+(v+k_{3}+1/2)\tilde{m}_{3}=0.$
Since $v=-k_{1}/(N-1)$ and $\tilde{m}_{2}=N-1$, one gets $(v+k_{3}+1/2)\tilde{m}_{3}=0$. Since $\tilde{m}_{3}>0$, we have $v=-k_{3}-1/2$, and so $w=v+k_{3}+1/2=0$, contrary to the fact $w\neq u$. ∎
Proof of Theorem 1.4.
(i) Suppose that $\mathfrak{X}$ is skew-symmetric. By Theorem 4.1 and Propositions 4.2, 4.3, $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues for some $i\in\{1,2,3\}$. Suppose that $\mathfrak{X}$ is not skew-symmetric. It follows that $\mathfrak{X}$ has exactly one pair nonsymmetric relations. Then the symmetrization of $\mathfrak{X}$ is a $3$-class association scheme. In view of Theorem 1.3, we only need to consider the case that the symmetrization of $\mathfrak{X}$ is not amorphic. By Lemma 4.4 and Propositions 4.64.8, $(X,R_{i}\cup R_{3})$ has $5$ distinct eigenvalues for some $i\in\{1,2,3\}$. Thus, (i) is valid.
(ii) is immediate from (i) and Theorem 2.6. ∎
Acknowledgements
Y. Yang is supported by NSFC (12101575, 52377162).
Data Availability Statement
No data was used for the research described in the article.